OpenStudy (razzanide):

For each function, y varies directly with x. Find each constant of variation. Then find the value of y when x=0.3. 13. y=2/3 when x=0.2 I know that you set it up as y=kx -> 2/3=k0.2 but I don't know how to solve it where the answer would be one I could work with? Because I have to then take that 'answer' and put it in the problem like so: y=-4x (and let 'answer'=x)

hartnn (hartnn):

find the value of k first! 2/3 = k* 0.2 correct :) find k from here.

OpenStudy (razzanide):

That's what I'm having trouble on? Like I'd have to either make the decimal into a fraction or vice versa, so that it would either come out as 0.6666...=k0.2 or 10/15=k3/15 and I'm not sure how to find k from either one of those without the problem answer becoming too weird to solve with?

hartnn (hartnn):

lets keep it in fraction only if we go for decimals, we need to do approximations so, 2/3 = k * 2/10 k= 2/3 times 10/2 = 10/3 keep k as 10/3 only

OpenStudy (razzanide):

So it'd be y=-(10/3)(3/10) for my final equation to solve?

hartnn (hartnn):

where did the negative sign come from ?

OpenStudy (razzanide):

oops, that was a typo

hartnn (hartnn):

yeah, without the negative sign, thats correct! just simplify :)

OpenStudy (razzanide):

Is my answer 1?

hartnn (hartnn):

thats correct :)

OpenStudy (razzanide):

Thank you so much!

hartnn (hartnn):

welcome ^_^